OpenStudy (anonymous):

insert 4 terms in geometric sequence between 3/8 and 128/81

OpenStudy (amistre64):

insert?

Parth (parthkohli):

lol, must be a homework question with 4 textboxes.

OpenStudy (amistre64):

\[\frac{3}{8},\frac{128}{81},-,-,-,-\] or \[\frac{3}{8},-,-,-,-,\frac{128}{81}\] ???

OpenStudy (anonymous):

the second one sorry it was meant to say between 3/8 and 128/81

OpenStudy (amistre64):

do you recall the general setup for a geometric seq? that might prove useful

OpenStudy (amistre64):

it can be done with means, but i find that technique a little obtuse :)

OpenStudy (amistre64):

given:\[g_1=\frac38~:~g_6=\frac{128}{81}\] and the general formula for a geo seq:\[g_n=g_1~r^{n-1}\] we can solve this for "r"\[\frac{g_n}{g_1}=r^{n-1}\]\[\left(\frac{g_n}{g_1}\right)^{\frac{1}{n-1}}=r\] use g6 to determine the r value to use in finding g2 g3 g4 and g5